OpenStudy (anonymous):

Solve x2 + 10x + 24 = 0 by completing the square

OpenStudy (anonymous):

this is what I have so far x2+10x+24=0 -24 -24 -24=x2+10x

OpenStudy (anonymous):

i guess you are in class 10 right?

OpenStudy (anonymous):

im using someone open study that was class 2012 which was my sister im in 9th doing early graduation course

OpenStudy (anonymous):

okay..you must go through the examples of your book that is going to help you the most right now.

OpenStudy (anonymous):

yes but I cant figure out if im using standard or general form? Do you even know how to help me on this question im not looking for the answer im looking to work through the question to figure out the answer

OpenStudy (mathstudent55):